OpenStudy (anonymous):

in solving bollean algebra problem ------ - - x=a.b+c+a.c+b here there is a whole bar on a.b+c and also a bar on c.there is another bar +a.c where c have also bar. plzzz help me

OpenStudy (saifoo.khan):

o.O

OpenStudy (saifoo.khan):

Lahori, right?

OpenStudy (anonymous):

\[\bar a + a = 1\] Is that right?

OpenStudy (anonymous):

and (c+a)(c+b)= c + ab Is that the type of logic we are talking about?

OpenStudy (anonymous):

\[\overline{a.b+c}+\overline{c}+\overline{a.\bar c}+b\] ?

OpenStudy (anonymous):

without the c in the middle I think

OpenStudy (anonymous):

\[\overline{a.b+c}+\overline{a.\bar c}+b\]

OpenStudy (anonymous):

This expression: not(a and b or not(c)) or b or a and not(c) Is equivalent to (see wolfram below): a or b or c Which needs to be shown with boolean algebra (see a later post made, which seems to refer to this question). http://www.wolframalpha.com/input/?i=not%28a+and+b+or+not%28c%29%29+or+b+or+a+and+not%28c%29

OpenStudy (anonymous):

I hate those equation because I'm ALWAYS making careless mistakes \[\overline{a.b+\bar c}+\overline{a.\bar c}+b = \overline{a.b}.c+\bar a+c+b=c.(\overline{a.b}+1)+\bar a + b=c.1 +\bar a +b=\bar a+b+c\]

OpenStudy (anonymous):

Is it \[\overline{a.b+\bar c}+\overline{a.\bar c}+b\] or \[\overline{a.b+\bar c}+a.\bar c+b\] ?

OpenStudy (anonymous):

I think it's the 2nd one you posted, and that you are supposed to show it is equivalent to A+B+C (see the next question that Muhammad posted). According to wolfram, the 2nd expression you posted IS equivalent to A+B+C .

OpenStudy (anonymous):

Yep, the second IS A + B + C Here is the simplification (at least my try) \[\overline{a.b+\bar c}+a.\bar c+b = \overline{a.b}.c+a.\bar c+b=(\bar a +\bar b).c+a.\bar c+b=\bar a.c+\bar b.c+a.\bar c +b\]\[=(\bar a.c+a.\bar c) +(\bar b.c+b)=a \oplus c +b +c = (a \oplus c +c) +b= a + c +b\]